Please derive from scratch! (Meaning, first and second laws for thermo or initial definitions of kinetics in terms of rate laws and rate expressions) 1. Here is a formula used to describe the separation process in gas chromatography: Inkvap + -pm-+ constant RT RT Which simplifies further for nonpolar molecules to: +constant RT RT Show how this final form is derived 2. Derive the Henderson Hasselbalch equation pH pKa+log ([base]/[acidl) First, start with a derivation of the equilibrium constant Ka and then consider a buffer where the degree of dissociation is small in either direction Derive the amphiprotic equation from mass and charge balance of a diprotic acid, again starting with the derivation of the equilibrium constants: 3.
Answer #1

1.

1st law

U = Q+ W

W= P \Delta V

Second law we get

S = Q/T

combing both we get

\Delta U = T dS - PdV

Also

\Delta H = T dS + V dP

\Delta H - T dS = \Delta G

-RT lnK = \Delta G

lnK = - \Delta H/ RT + \Delta S/R

which is our eqn

2.

From dissociateion of acid in water we get

K_{a} = [H+][A-]/ [HA]

logK_{a} = log([H+][A-]/ [HA])

-pKa= log([H+]) + log([A-]/ [HA])

-pKa=-pH + log([A-]/ [HA])

pKa=pH - log([A-]/ [HA])

is the henderson hasselbach eqn